OpenStudy (anonymous):

need to know which number is irrational and rational and then change to to decimals (to the nearest hundredth), 1 fourth, o.3 with the line on top of the 3, squareroot 3, 15 over 8, and negative 1.247

Directrix (directrix):

A rational number is a number that can be written as the ratio of two integers. 1 fourth = 1/4 Rational o.3 with the line on top of the 3 = 1/3 Rational squareroot 3, Irrational 15 over 8 = 15/8, Rational - 1.247 = =1247/1000, Rational